# Cache Me If You Must: Adaptive Key-Value Quantization for Large Language Models

### v1

@@ 1 Introduction
    We propose AQUA-KV — a novel compression framework that exploits inter- and intra-layer dependencies to improve quantization accuracy. AQUA-KV works in one-shot, based on a lightweight calibration procedure, and shows competitive size-accuracy trade-offs. Additionally, it is compatible with arbitrary quantization techniques, and can be combined with additional compression, such as pruning.
@@ 3.1 Analysis of Inter-Layer Dependencies
To better quantify these errors, note that 1- and 2-bit quantizers usually explains 0.75 and 0.89 variance, respectively. Intuitively, if a probe can predict the $L$-th block keys with the same relative error as the 1- or 2-bit quantizer, it means that using the probe can “save” approximately this many bits per value when quantizing residuals. While this is not a strict guarantee, we found that it holds well for real-world LLMs, as can be seen in Sections 4.1 and 4.2.
@@ 3.3 Implementation Details
Relation between AQUA-KV and Attention Sinks. It is well-known that modern LLMs tend to form attention “sinks” — tokens that have extremely high attention scores while being semantically unimportant (Xiao et al., 2024). Several prior works in KV compression propose special treatment for such attention sinks, such as keeping them in higher precision (Hooper et al., 2024) or introducing synthetic sinks (Xiao et al., 2024; Chen et al., 2025).
@@ 4 Experiments
When evaluating perplexity, we adopt the same approach as in prior quantization works (Frantar et al., 2022; Lin et al., 2023; Egiazarian et al., 2024), with one exception: instead of processing sequences (of length 8192) in parallel, we encode them auto-regressively and maintain the (compressed) Key-Value cache as during as during inference. This results in the same perplexity for non-compressed KV cache, but allows us to properly account for the effect of recent token buffers and attention sinks during KV quantization, as described in Section 3.3. We evaluate perplexity on base (non-instruct) models since they have better .
@@ 4.1 Detailed Evaluation & Ablation Analysis
KVSharer (1 pair) 15.43 7.45 36.81
KVSharer (4 pairs) 13.71 9.60 29.82
Predictor Architecture
@@
1st Layer Keys & Values
@@ 4.3 Compatibility with Pruning
Bits 3B 8B
H2O only 16 38.82 41.42
HIGGS 2.02 37.02 40.53
AQUA-KV 2.09 38.43 40.88
@@
We ran 8B model on a single A100 GPU and the 70B on 2xA100 in sequential mode, generating a single sequence of 8192 tokens with 2-bit HIGGS. For the 8B model, generating with HIGGS quantizer runs at 24.02 tokens/second and AQUA-KV runs at 23.31. For 70B, HIGGS runs at 5.91 tokens/s and AQUA-KV at 5.76 tokens/s, for the average difference of about 3%. However, this slowdown can be outweighed by the substantially better accuracy of AQUA-KV.
@@ Appendix C On LongBench Evaluation on non-Instruct models
The tables 5-9 summarize the average LongBench performance across all 14 tasks for each model, followed by the WikiText-2 perplexity scores. Additionally, detailed per-task results are provided for a deeper understanding of how different models and quantization settings affect specific NLP capabilities.

### v2

@@ 1 Introduction
    We propose AQUA-KV — a novel compression framework that exploits inter- and intra-layer dependencies to improve quantization accuracy. AQUA-KV works in one shot, based on a lightweight calibration procedure, and shows competitive size-accuracy trade-offs. Additionally, it is compatible with arbitrary quantization techniques, and can be combined with additional compression, such as pruning.
@@ 2.1 KV-Cache Compression
Linearity between adjacent layers.
Prior work (Razzhigaev et al., 2024) has shown that there is a almost linear relationship between activations in sequential layers in transformer language models due to the presence of skip connections. We leverage this property for the design of KV-cache predictors.
@@ 3.1 Analysis of Inter-Layer Dependencies
We report these errors in Figure 3. For comparison, 1-bit and 2-bit quantizers usually explain 0.75 and 0.89 variance, respectively. Intuitively, if a probe can predict the $L$-th block keys with the same relative error as the 1- or 2-bit quantizer, it means that using the probe can “save” approximately this many bits per value when quantizing residuals. While this is not a strict guarantee, we found that it holds well for real-world LLMs, as can be seen in Sections 4.1 and 4.2.
@@ 3.3 Implementation Details
Relation between AQUA-KV and Attention Sinks. It is well-known that modern LLMs tend to form attention “sinks” — tokens that have extremely high attention scores while being semantically unimportant (Xiao et al., 2023). Several prior works in KV compression propose special treatment for such attention sinks, such as keeping them in higher precision (Hooper et al., 2024) or introducing synthetic sinks (Xiao et al., 2023; Chen et al., 2025).
@@ 4 Experiments
When evaluating perplexity, we adopt the same approach as in prior quantization works (Frantar et al., 2022; Lin et al., 2023; Egiazarian et al., 2024), with one exception: instead of processing sequences (of length 8192) in parallel, we encode them auto-regressively and maintain the (compressed) Key-Value cache as during as during inference. This results in the same perplexity for non-compressed KV cache, but allows us to properly account for the effect of recent token buffers and attention sinks during KV quantization, as described in Section 3.3. Here, we use base (non-instruct) models since they have better perplexity.
@@ 4.1 Detailed Evaluation & Ablation Analysis
AQUA-KV Predictor Architecture
@@
AQUA-KV 1st Layer Keys & Values
@@ 4.3 Compatibility with Pruning
Bits Saved 3B 8B
H2O only 16 5.0$\times$ 38.82 41.42
H2O + HIGGS 2.02 39.6$\times$ 37.02 40.53
H2O + AQUA-KV 2.09 38.3$\times$ 38.43 40.88
@@
We ran 8B model on a single A100 GPU and the 70B on 2xA100 in sequential mode, generating a single sequence of 8192 tokens with 2-bit HIGGS. For the 8B model, generating with HIGGS quantizer runs at 24.02 tokens/second and AQUA-KV runs at 23.31. For 70B, HIGGS runs at 5.91 tokens/s and AQUA-KV at 5.76 tokens/s, for the average difference of about 3%. However, this slowdown can be outweighed by the substantially better accuracy of AQUA-KV. Note that these results were obtained using unoptimized PyTorch and can likely be improved with specialized libraries.
@@ Appendix C On LongBench Evaluation on non-Instruct models
The Tables 6-10 summarize the average LongBench performance across all 14 tasks for each model, followed by the WikiText-2 perplexity scores. Additionally, detailed per-task results are provided for a deeper understanding of how different models and quantization settings affect specific NLP capabilities.